A bottle full of mercury has a mass of 720 grams,when full of water its mass is 90 grams.Calculate the volume of the empty density bottle,the mass of the bottle and the density of the mercury?
1
Expert's answer
2020-04-30T10:42:27-0400

Mass of bottle is 40g.

"RD=\frac{720-40}{90-40}=13.6"

"\rho_{Hg}=13600\frac{kg}{m^3}"

"0.09-0.04=(1000)(9.8)V"

"V=5.1\cdot10^{-6}m^3"